### v2

# Thouless approach in transport in integrable and perturbed easy-axis Heisenberg chains

Abstract
We study transport in spin chains employing the Thouless approach based on the level sensitivity to the boundary conditions, $R$. Although spin transport in the integrable easy-axis XXZ model is diffusive, corresponding $R$ is much closer to ballistic chains than to chaotic diffusive systems. In the case of the grand canonical ensemble, this observation can be rigorously justified, while in the case of the canonical ensemble it can be demonstrated by numerical calculations. Integrability breaking perturbation (IBP) strongly reduces $R$ which reveals a pronounced minimum at the crossover from anomalous diffusive to normal dissipative transport. This minimum coincides with the onset of the universality of the random matrix theory. Results for various IBP suggest a discontinuous jump of the spin conductivity in the thermodynamic limit, and moreover that its value is proportional to the strength of the IBP.
Introduction. The properties of integrable quantum lattice models, the prominent example being the one-dimensional, anisotropic Heisenberg XXZ spin-$1/2$ chain, have in last decades reached higher level of understanding (for recent review see e.g. [1]) through the extended concepts of local [2] and quasilocal conserved quantities (CQ) [3, 4], via the analytical results emerging from the generalized hydrodynamics (GHD) [5, 6, 7, 8, 9] and with the application of various powerful numerical methods [10, 11, 12, 13, 14, 15]. While the standard argument is that mastering integrable models will be the crucial step to the proper description of closely related perturbed models, the advance in this direction is so far modest. In the case of the Heisenberg-type spin chains it appears evident that the integrability-breaking perturbations (IBP) suppress the finite-temperature ballistic transport, characterized by the spin stiffness $D(T>0)>0$, turning it into the normal dissipative one [16, 17, 18, 19, 20]. Still, analytical approaches and tools dealing with such perturbed systems are so far quite restricted [21, 22, 23] .
The anisotropic Heisenberg chain in the easy-axis ($\Delta>1$) regime represents a quite different challenge, with vanishing spin conductivity (and diffusion) at temperature $T\to 0$, as well as vanishing stiffness $D(T>0)=0$, but revealing finite spin diffusion constant ${\cal D}$ even at high $T$ [11, 24, 13, 25]. The diffusion is still anomalous in origin [26, 27, 28], and the transport is dissipationless [15]. Taking into account open questions in this problem, the introduction of IBP leads to even wider range of scenarios. While there appears general consensus that in large enough systems (system size $L\to\infty$) the spin transport becomes normal (i.e. dissipative), its variation with IBP strength $g$ remains controversial. While some recent results indicate on the continuous variation of ${\cal D}$ with $g$ [29], there is also evidence for its discontinuity at $g\to 0$ [15, 30]. Moreover, introducing in the system a field $F>0$ [12] even favors the vanishing of the spin conductivity with the vanishing $F\to 0$.
In this Letter we use an alternative approach to the spin transport both in integrable and perturbed XXZ chains at $\Delta>1$. Since these spin chains are equivalent to the models of interacting fermions, one can introduce a nonzero flux [31] that probes also system’s response to changing of the boundary conditions. In particular, we employ the Thouless level sensitivity $R$. It was originally introduced into single-particle models to distinguish between conductors with $R\gg 1$ and insulators (localized systems) with $R\to 0$ [32] and has recently been applied also to interacting disordered systems [33].
The level sensitivity distinguishes between ballistic and normal diffusive transport. In the former case it grows (up to logarithmic corrections) as the inverse of the mean level spacing, $R\propto 1/\Delta\epsilon$ whereas in generic diffusive systems this dependence is much weaker, namely $R\propto 1/\sqrt{\Delta\epsilon}$. We show that the diffusive transport in the easy-axis integrable XXZ chain ($\Delta>1$) is anomalous in this respect. Namely, diffusive integrable chains reveal $R\propto 1/\Delta\epsilon$ in the grandcanonical ensemble (GCE), as well as in the canonical ensemble (CE) at total $S^{z}_{tot}=0$. Therefore, the level sensitivity of diffusive integrable system differs from $R$ in the ballistic systems only by logarithmic corrections, $\log(\Delta\epsilon)$, and is much larger than in generic diffusive models. Large values of the level sensitivity in GCE, $R_{\text{GCE}}\gg 1$, are explained by strict lower bound.
We also observe that level sensitivity in CE, $R_{\text{CE}}$, obtained for finite systems exhibits a nonmonotonic dependence on the strength of IBP, $g$. The minimum at $g=g^{*}(L)$ coincides with the crossover where the finite-size system starts to comply with the random matrix theory (RMT) [34, 35]. Therefore, the minimum of $R$ allows us to delimit two regimes: $g>g^{*}(L)$ when the system behaves as normal dissipative and $g<g^{*}(L)$ when the system properties are still dominated by finite-size effects. Taking into account the restriction, $g>g^{*}(L)$, one can then follow in a controlled way the variation with $g$ of the d.c. spin conductivity in CE, $\sigma_{\text{CE}}$. Our results for two classes of IBP seem to support a universal limit, i.e., vanishing of $\sigma_{\text{CE}}(g\to 0)$ in the thermodynamic limit.
XXZ Model with flux. We consider the XXZ spin ($s=1/2$) chain of length $L$ with periodic boundary conditions (PBC), introducing also a flux, $\varphi$,
$H=J\sum_{l}\left[\frac{1}{2}(\mathrm{e}^{-i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c.})+\Delta S^{z}_{l+1}S^{z}_{l}\right]+gH^{\prime}.$ (1)
We consider the easy-axis regime with $\Delta>1$ and set $J=1$. The spin model can be mapped on the model of interacting spinless fermions, where $\varphi L$ represents a magnetic flux through the ring [31, 32, 15]. We analyze two forms of IBP. In the main text, we discuss the next-nearest neighbor (nnn) interaction with $g=\Delta_{2}$ and $H^{\prime}=\sum_{l}S^{z}_{l+2}S^{z}_{l}$, whereas in the Supplemental Material [36] we study the nnn exchange with $g=J_{2}$ and $H^{\prime}=\sum_{l}[\mathrm{e}^{-2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}]$ which maps to a nnn hopping in the fermionic chain.
Figure 1: (a) GCE: bound from Eq. (5) (lines), compared with ED results (points) for the diagonal matrix elements, $\overline{j^{2}_{nn}}$. (b) CE: $\overline{j^{2}_{nn}}$ obtained via ED for the $S^{z}_{tot}=0$ and various $\Delta=0.5-4$. The horizontal line is the GCE bound, Eq. (5), that holds for $\Delta\geq 1$ and any $L$.
Level sensitivity and spin transport. The gauge transformation $S^{\pm}_{l}\to e^{\pm il\varphi}S^{\pm}_{l}$, $S^{z}_{l}\to S^{z}_{l}$ applied to the Hamiltonian (1) eliminates the flux dependence on all sites except the boundary term that takes the form $J/2(S^{+}_{1}S^{-}_{L}e^{-i\varphi L}+{\rm H.c.})$. Then, the energy spectra, $H|n\rangle=\epsilon_{n}(\varphi)|n\rangle$, obtained for $\varphi=0$ and $\varphi=\delta\varphi=\pi/L$ correspond, respectively, to periodic and antiperiodic ($J\to-J$) boundary conditions. Consequently, the level sensitivity to the boundary conditions is related to the derivative $\epsilon^{\prime}_{n}(\varphi)$ that can be further simplified using an identity for the spin-current operator, $j={\rm d}H/{\rm d}\varphi$. Following the reasoning from Refs. [31, 37] one finds
$\epsilon^{\prime}_{n}(\varphi)=\frac{{\rm d}}{{\rm d}\varphi}\langle n|H|n\rangle=\langle n|j|n\rangle+\epsilon_{n}\frac{{\rm d}}{{\rm d}\varphi}\langle n|n\rangle=\langle n|j|n\rangle.$ (2)
The level sensitivity introduced by Thouless in Ref. [32] is determined by the matrix elements of the spin-current operator, $j_{mn}=\langle n|j|m\rangle$,
$R=\delta\varphi\sqrt{\overline{(\epsilon^{\prime}_{n}(\varphi))^{2}}}/\Delta\epsilon=\delta\varphi\;\sqrt{\overline{j^{2}_{nn}}}/\Delta\epsilon,\quad\delta\varphi=\frac{\pi}{L}.$ (3)
Here, overline means averaging over the eigenstates (in the relevant part of the spectrum) and $\Delta\epsilon$ is the average level spacing. We note that integrable XXZ model with $\Delta<1$ exhibits a ballistic transport. It is quantified by the spin stiffness (Drude weight) [2], $D=\overline{j^{2}_{nn}}/L>0$ that is finite in the thermodynamics limit $L\to\infty$. Consequently, in ballistic models one obtains $R=\pi\sqrt{D}/(\sqrt{L}\Delta\epsilon)$.
In the case of the easy-axis regime, $\Delta>1$, the ballistic component vanishes in GCE and in CE with $S^{z}_{tot}=0$. In both cases one finds $D=0$ for $L\to\infty$. However in the GCE, one may still find a simple bound on the level sensitivity. To this end we employ the Mazur bound [38, 2], $\overline{j^{2}_{nn}}\geq\langle jQ\rangle^{2}/\langle Q^{2}\rangle$, where $Q$ is a conserved operator, $\langle...\rangle={\rm Tr}(...)/N_{st}$ and $N_{st}$ is the Hilbert-space dimension. We choose $Q$ as a product of two simplest local conserved operators, $Q=Q_{1}Q_{3}$, where $Q_{1}=\sum_{l}S^{z}_{l}$ is the total magnetization and $Q_{3}$ is the energy current,
$Q_{3}$ $=$ $\,\frac{iJ}{2}\sum_{l}\left(JS_{l+1}^{-}S_{l}^{z}S_{l-1}^{+}+\Delta S_{l+1}^{z}S_{l}^{+}S_{l-1}^{-}\right.$ (4)
$+\left.\Delta S_{l+1}^{+}S_{l}^{-}S_{l-1}^{z}+\mathrm{H.c.}\right).$
Direct calculations in GCE give a bound
$\overline{j^{2}_{nn}}\geq\frac{\langle jQ_{1}Q_{3}\rangle^{2}}{\langle Q^{2}_{1}Q^{2}_{3}\rangle}=\frac{J^{2}\Delta^{2}L}{2J^{2}(L-2)+4L\Delta^{2}},$ (5)
We present in Fig. 1(a) the GCE bound for different $\Delta\geq 1$ and compare it with the numerically evaluated value $\overline{j^{2}_{nn}}$ obtained in GCE for systems with $L\leq 18$. One can notice, that the lower bound even quantitatively approaches the numerical value at $\Delta\gg 1$. In the easy-axis regime, $\Delta>1$, that is of interest here, one finds from Eq. (5) $\overline{j^{2}_{nn}}\geq 1/6$ and $R\geq(0.4\pi)/(L\Delta\epsilon)$. In the GCE, the level sensitivity in ballistic and diffusive regimes of the integrable XXZ chains differ only by a factor $1/\sqrt{L}\sim{\cal O}[\log(\Delta\epsilon)]$.
The above bound is inapplicable in the CE with $S^{z}_{tot}=0$ because $Q_{1}=0$. Fig. 1(b) shows finite-size results for $\overline{j^{2}_{nn}}$. Here and further on, they are obtained using exact diagonalization (ED). For $L\leq 22$, we carry out full diagonalization for sectors with the wavevectors $0<q<\pi$ and use half of the states from the middle of the spectra. For the largest $L=24$ we employ the shift-invert ED and randomly sample over $N\sim 150$ eigenstates from the middle of the spectra, in each sector with $0<q<\pi$ (see Supplemental Material [36] for details).
One observes that $\overline{j^{2}_{nn}}$ is well below the GCE bound and that the difference between GCE and CE results increases with $L$ and $\Delta$. In order to elucidate such $\Delta$-dependence, we have numerically studied also the ’folded’ XXZ model [39, 40, 30] that represents $\Delta\to\infty$ limit of the XXZ chains, $H_{f}=(J/2)\sum_{l}P_{l}\left[\mathrm{e}^{-i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c}\right]$ with the projection operator $P_{l}=(S^{z}_{l+2}+S^{z}_{l-1})^{2}$ that introduces conservation of the number of domain walls. In this model we find (so far numerically) strictly $\overline{j^{2}_{nn}}=0$ for any $L$. Consequently, one gets $R_{\text{CE}}=0$ while $R_{\text{GCE}}>0$. The latter follows from the strict bound in Eq. (5) and originates from the ballistic transport in all sectors with $S^{z}_{tot}\neq 0$. Analytical support for vanishing $R_{\text{CE}}$ is discussed in the Supplemental Material. [36] Following results in Fig. 1(b) we present in Fig. 2(a) the dependence of $R_{\text{CE}}$ on the average level spacing $\Delta\epsilon$. Results show that for finite $\Delta<\infty$, the anomalous dependence $R\propto 1/\Delta\epsilon$ indeed holds true, being thus close to ballistic systems rather than to generic diffusive ones.
The properties of $R$ in chaotic models are related to the d.c. spin conductivity $\tilde{\sigma}$ via the RMT relations [41, 42, 43]. Since at high $T\gg J$ one gets $\tilde{\sigma}\propto T^{-1}$, we consider here a rescaled $\sigma=T\tilde{\sigma}$, defined as the (Kubo) linear-response function
$\sigma=\lim_{\omega\to 0}\frac{\pi}{LN_{st}}\sum_{n\neq m}|j_{mn}|^{2}\delta(\omega-\epsilon_{m}+\epsilon_{n})\simeq\frac{\pi\overline{|j_{mn}|^{2}}}{L\Delta\epsilon},$ (6)
where $\overline{|j_{mn}|^{2}}$ are averaged matrix elements for small $|\epsilon_{m}-\epsilon_{n}|$. Upon introducing a flux $\varphi\neq 0$, the degeneracy of energy levels persists only in momentum sectors $q=0$ and $q=\pi$. All other sectors obey the RMT relations [43], which in our system follow the Gaussian orthogonal ensemble (GOE) universality, requiring $Y=\overline{|j_{mn}|^{2}}/\overline{j_{nn}^{2}}=1/2$. Therefore, RMT establishes a link between the level sensitivity, Eq. (3) and the spin conductivity, Eq. (6),
$R=\sqrt{\frac{2\pi\sigma}{L\Delta\epsilon}}.$ (7)
Eq. (7) means that the level sensitivity in chaotic systems scales with the system size as $R\propto 1/\sqrt{L\Delta\epsilon}$. It is a general (i.e. model independent) property that follows solely from RMT and the finiteness of $\sigma$. In the case of integrable systems, $R$ should follow a simple bound $R\geq c/(L^{m+1}\Delta\epsilon)$ where $c$ is a constant of order one. Independently of the studied model, a sufficient condition for the latter to hold is that the current operator has a non-vanishing projection on products of a finite number of LIOMs, $\langle jQ_{1}Q_{2}\ldots Q_{m}\rangle\neq 0$, see the discussion in the Supplemental Material [36]. We stress that the latter projection implies ballistic transport only for $m=1$. We also recall that the level spacing, $\Delta\epsilon$, decays exponentially with $L$. Consequently, introducing of IBP should in general be accompanied by a substantial drop of the level sensitivity from $R\propto 1/\Delta\epsilon$ in the integrable chain (either ballistic or diffusive) to $R\propto 1/\sqrt{\Delta\epsilon}$ in the chaotic system. All other terms which enter $R$ can be bounded by a power-law function $\sim 1/L^{m+1}$ so that they appear only as logarithmic corrections, $\sim\log(\Delta\epsilon)$.
Figure 2: Results for CE with $S^{z}_{tot}=0$ from full ED (filled symbols) and shift-invert method (open symbols). (a) Scaled level sensitivity $LR_{\text{CE}}$ vs $1/\Delta\epsilon$, where $\Delta\epsilon$ is average level spacing in sectors with fixed momenta. Dashed lines are guidelines showing $1/\Delta\epsilon$ and $1/\sqrt{\Delta\epsilon}$ dependence, respectively. (b)-(d) $R_{\text{CE}}$ vs. IBP strength $\Delta_{2}$ for $\Delta=1.5,2.0,3.0$. Regions with $R<1$ are shaded, whereas black triangles correspond to minima of $R(\Delta_{2})$, approximated from a quadratic interpolation of three subsequent data points.
Figure 3: RMT indicators for the perturbed model in canonical sector. a) Gap ratio $r$, inset: ratio $Y$ of off-diagonal and diagonal matrix elements. b) Kurtosis $K$ (defined in text).
All result henceforth apply to CE. Figure 2(a) confirms that $R_{\text{CE}}$ in the integrable XXZ model shows the dependence $R_{\text{CE}}\propto 1/\Delta\epsilon$ for all finite $\Delta$. Comparing this result with $R_{\text{CE}}=0$ in the ’folded’ model, one finds that the limits $L\to\infty$ and $\Delta\to\infty$ do not commute. The limit $\Delta\to\infty$ is taken first in the ’folded’ model. Fig. 2(a) confirms also that substantial IBP ($g=\Delta_{2}$) causes qualitative changes of $R_{\text{CE}}$, as discussed in the preceding paragraphs. More systematic dependence of $R_{\text{CE}}$ on the IBP strength $\Delta_{2}$ is presented in Figs. 2(b)- 2(d) for different $\Delta>1$. One observes that $R_{\text{CE}}$ develops a clear minimum at finite $\Delta_{2}$ that drifts towards weaker IBP when $L$ increases. It should be noted, that we are interested in the regime $R_{\text{CE}}\gtrsim 1$. For $R_{\text{CE}}\lesssim 1$ (shaded areas in Fig. 2) the system is too small, i.e. the level spacing is too large, so that changing of the flux causes neither level crossings nor avoided crossings (within the same $S^{z}_{tot}$ and momentum sector).
Figure 4: Spin conductivity in canonical sector, $\sigma_{\text{CE}}$, for: (a) $\Delta=1.5$ and (b) $\Delta=2.0$. Results obtained from diagonal matrix elements via full ED ($L=18-22$) and shift-invert ($L=24$) are presented in the regime $\Delta_{2}>\Delta_{2}^{*}(L)$, i.e., above the crossover to the GOE universality. Presented are also MCLM results for $L=28$.
In order to explain the origin of this minimum, we numerically study the standard indicators of RMT. The main panel in Fig. 3(a) shows the gap ratio $r=\overline{r}_{n}$ with $r_{n}=\mathrm{min}[\Delta_{n},\Delta_{n+1}]/\mathrm{max}[\Delta_{n},\Delta_{n+1}]$, where $\Delta_{n}=\epsilon_{n+1}-\epsilon_{n}$. The inset shows the ratio of the off-diagonal and the diagonal matrix elements, $Y=\overline{|j_{mn}|^{2}}/\overline{|j_{nn}|^{2}}$. Both results confirm that sufficiently strong IBP introduces the GOE universality [44, 45] with $Y=0.5$ and $r\simeq 0.53$. Fig. 3(b) shows the kurtosis $K=\overline{j_{nn}^{4}}/\left(\overline{j_{nn}^{2}}\right)^{2}$. At large enough $\Delta_{2}$, $K$ agrees with the kurtosis of the Gaussian distribution with $K=3$. Comparing results in Figs. 2 and 3 one finds that minimum in $R_{\text{CE}}$ corresponds to the crossover to the normal RMT behavior. In other words, the minimum reveals the threshold strength of IBP, $g^{*}(L)$, when the scattering becomes strong enough to introduce the RMT universality.
In systems which satisfy RMT, the spin conductivity, $\sigma$, can be obtained also from the diagonal matrix elements, $\overline{|j_{nn}|^{2}}$, via Eqs. (3) and (7). Fig. 4 shows such results for the spin conductivity in CE, $\sigma_{\text{CE}}$. As before we use $g=\Delta_{2}$ as IBP. In Fig. 4 we present only numerical results which are beyond the crossover to the GOE universality, i.e. obtained for $g>g^{*}(L)$. The minima in $R_{\text{CE}}$ in Figs. 2(b)- 2(d) are here visible as the starting points of curves obtained for various L. In Fig. 4 we show also results obtained from the microcanonical Lanczos method (MCLM) [46, 47]. This method allows to evaluate full dynamical $\sigma_{\text{CE}}(\omega)$, and consequently also $\sigma_{\text{CE}}$ as in Eq. (6), for larger systems, but with smaller frequency resolution $\delta\omega$. Here, we present MCLM results for $L=28$ with $\delta\omega\lesssim 10^{-3}$ [15], which we estimate to be meaningful at least down to $g=\Delta_{2}\sim 0.15$.
We note that results for $\sigma_{\text{CE}}$ obtained from various methods are consistent provided that $g>g^{*}(L)$. One also observes that $\sigma_{\text{CE}}$ strongly depends on $\Delta$ resembling the analogous dependence of $R_{\text{CE}}$ on $\Delta$ shown Fig. 2(a). Apparently for $g>g^{*}(L)$, the spin conductivity follows the relation $\sigma_{\text{CE}}\propto|g|$. Here we can make a link with the diffusion constant, ${\cal D}$. In generic dissipative system it is related to $\sigma$ via the Einstein relation, that at large $T$ reads ${\cal D}=4\sigma$. The diffusion has been studied extensively in the integrable case and evaluated mostly in the GCE, both analytically within the GHD [25] and numerically via various methods [26, 11, 14, 13, 15] leading to values ${\cal D}_{\text{GCE}}>0.4$ in the regime $\Delta>1$ Evidently, $\sigma_{\text{CE}}$ shown in Fig. 4 for perturbation $g>g^{*}(L)$ is well below the latter. This discrepancy suggests a discontinuous change of ${\cal D}$ upon the introduction of IBP in agreement with Refs. [30, 15].
Conclusions and Discussion We have studied sensitivity of the energy levels to the flux (or equivalently twisted boundary conditions), $R$, and related transport quantities, within the XXZ chain in the easy-axis regime. Although the transport in the integrable model is diffusive, obtained $R$ is very close to one in integrable ballistic systems and, up to logarithmic corrections, exhibits the same dependence on the level spacing, $R\propto 1/\Delta\epsilon$. It strongly contrasts with chaotic diffusive systems where $R\propto 1/\sqrt{\Delta\epsilon}$. In the case of GCE , we have found for integrable chain a strict bound $R_{\text{GCE}}\geq 0.4\pi/(L\Delta\epsilon)$ that holds for arbitrary anisotropy $\Delta\geq 1$ and all system sizes, $L$. In the CE and for finite anisotropy $\Delta<\infty$, the level sensitivity still varies as $R_{\text{CE}}\propto 1/\Delta\epsilon$, however $R_{\text{CE}}$ is much smaller than $R_{\text{GCE}}$. The difference between $R_{\text{GCE}}$ and $R_{\text{CE}}$ grows with the systems sizes and anisotropy. In the limiting case $\Delta\to\infty$ we have found that $R_{\text{CE}}$ strictly vanishes for all system sizes whereas $R_{\text{GCE}}$ is bounded from below. In the Supplemental Material [36], we discuss results for nonzero magnetization sectors.
The qualitative difference between the level sensitivity in integrable and chaotic systems implies that a perturbation of strength $g$ must cause a substantial reduction of $R$. In finite systems, $R_{\text{CE}}$ is minimal at the crossover between integrable and RMT regimes. The minimum at $g=g^{*}(L)$ shifts towards weaker perturbations with increasing L.
In the chaotic regime, RMT establishes the Thouless-like relation between $R$ and the d.c. spin conductivity, $\sigma$. Using this relation we have numerically studied $\sigma$ in CE and found for $g>g^{*}(L)$ that $\sigma_{CE}\propto|g|$. For weak perturbations, $\sigma_{CE}$ becomes much smaller than $\sigma_{GCE}$ estimated from the diffusion constant for integrable systems. Assuming that the Einstein relation and the equivalence between CE and GCE hold true for all model parameters, our results support a discontinuous variation of $\sigma_{GCE}=\sigma_{CE}$ with IBP. It should be acknowledged that the limit of a weak perturbation is highly nontrivial since the difference between CE and GCE results goes as, $1/(g^{2}L)$ [15]. Consequently, the convergence of results in both ensembles is hard to establish numerically.
Acknowledgments. We acknowledges fruitful discussions with L. Zadnik and R. Steinigeweg. M.M. acknowledges support by the National Science Centre (NCN), Poland via project 2020/37/B/ST3/00020. J.P. acknowledges support by the National Science Centre (NCN), Poland via project 2023/49/N/ST3/01033. P. P. acknowledges support of the Slovenian Research Agency via the program P1-0044.
Supplemental Material:
Thouless approach in transport in integrable and perturbed easy-axis Heisenberg chains
J. Pawlowski 1, M. Mierzejewski1, and P. Prelovšek2
1 Institute of Theoretical Physics, Faculty of Fundamental Problems of Technology, Wrocław University of Science and Technology, 50-370 Wrocław, Poland
2J. Stefan Institute, SI-1000 Ljubljana, Slovenia
In the Supplemental Material we present additional results with next-nearest-neighbor exchange as the perturbation, details of shift-invert calculations, and extra considerations of the ’folded’ XXZ model, which can serve as the toy model for the behavior of the easy-axis $\Delta>1$ regime. We also show an analytical argument bounding the diagonal matrix elements of the spin current. Finally, we discuss the level sensitivity for nonzero magnetization sectors.
S0.1 Next-nearest-neighbor exchange as the perturbation: results
In addition to the next-nearest-neighbor (nnn) interaction $g=\Delta_{2}$ taken as the integrability-breaking perturbation (IBP) we present analogous results also for the the nnn exchange where $g=J_{2}$ and
$H^{\prime}=\sum_{l}[\mathrm{e}^{-2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}].$ (S1)
We note that such term maps to a nnn hopping in the fermionic chain.
Figure S1: Level sensitivity $R_{CE}$ vs. perturbation strength $J_{2}$ within sector $S^{z}_{tot}=0$, as calculated via full ED in systems with $L=18-22$ sites, for: (a) $\Delta=1.5$, (b) $\Delta=2.0$, and (c) $\Delta=3.0$. Black triangles correspond to minima of $R(\Delta_{2})$, approximated from a quadratic interpolation of three subsequent data points.
We first present in Fig. S1 results for the level sensitivity $R$ varying the IBP strength $J_{2}$, as calculated via full ED in systems with $L=18-22$. Again, substantial IBP ($J_{2}\gg 0$) causes qualitative changes of $R$, inducing a minimum of $R(J_{2})$ indicating the crossover into the regime $J_{2}>J_{2}^{*}(L)$ with RMT universality. As in the main text, also here we restrict to the regime $R>1$ where the interpretation of results in terms of normal transport is meaningful. While the behavior is quite similar to the case of $g=\Delta_{2}$, presented in the main text, it should be noticed that the dependence of $R$ on $\Delta$ in Fig. S1 is less pronounced than for $\Delta_{2}$. Larger values of $R$ in the present case allow us also to reach $R>1$ also for $\Delta=3$. The qualitative difference emerges at $\Delta\to\infty$ where in the folded model with $\Delta_{2}\neq 0$ one obtains $R=0$, while this is not the case for $g=J_{2}$ which introduced $R>0$ even in the folded model.
Figure S2: RMT indicators for the model with $J_{2}$ perturbation. a) Gap ratio $r$, inset: ratio $Y$ of off-diagonal matrix elements versus the diagonal ones. b) Kurtosis $K$ (defined in the text).
We present in Fig. S2 results for standard indicators of RMT: in Fig. S2(a) the gap ratio $r=\overline{r}_{n}$ with $r_{n}=\mathrm{min}[\Delta_{n},\Delta_{n+1}]/\mathrm{max}[\Delta_{n},\Delta_{n+1}]$, in the inset of Fig. S2(a) the ratio $Y=\overline{|j_{mn}|^{2}}/\overline{|j_{nn}|^{2}}$, and finally in Fig. S2(b) the kurtosis $K=\overline{j_{nn}^{4}}/\left(\overline{j_{nn}^{2}}\right)^{2}$. All results again confirm that sufficiently strong IBP $g>g^{*}(L)$ introduces the GOE universality with $Y=0.5$, $r\simeq 0.53$, and $K\simeq 3$.
Figure S3: Spin conductivity $\sigma_{CE}$ for (a) $\Delta=1.5$ and (b) $\Delta=2.0$. Points denote results obtained from diagonal matrix elements via full ED ($L=18-22$), presented in the regime $J_{2}>J_{2}^{*}(L)$, i.e., above the crossover to the GOE universality. Presented are also MCLM results for $L=24$ and $L=28$.
Finally, we show results for $\sigma_{CE}$ beyond the crossover to the GOE universality $g>g^{*}(L)$ for $g=J_{2}$, here obtained via ED with up to $L=22$. Again, results for $\sigma_{CE}$ become quantitatively consistent (weakly $L$-dependent) provided that $g>g^{*}(L)$. Extrapolating the crossover $g^{*}(L)$ to larger $L$, we can compare then results to those obtained via MCLM for larger $L=24$ and $L=28$, which then apparently leads to RMT-meaningful results down to $g=J_{2}\sim 0.2$ as presented in Fig. S3. The conclusion are analogous to one presented for $g=\Delta_{2}$. The difference again is that the dependence of $\sigma_{CE}$ on $\Delta$ for $g=J_{2}$ is weaker than for $g=\Delta_{2}$. Namely, $\sigma_{CE}$ at $J_{2}>0$ remains substantial also in the limit $\Delta\to\infty$ , i.e., within the folded model. Otherwise, for $J_{2}\ll 0.5$ the values of $\sigma_{CE}$ are well below the integrable GCE value $\sigma_{GCE}>0.1$ indicating on the discontinuous jump at $g\to 0$. Moreover, the variation of $\sigma_{CE}$ with $g$ in Fig. S3 reveals $\sigma_{CE}\propto|g|$ for $g>g^{*}(L)$. However, as in the main text we stress that the regime of weak IBP should be taken with care and requires further attention.
S0.2 Numerical methods
We evaluate the diagonal matrix elements of the spin current for $L=24$ using a custom implementation of the shift-invert method [49], employing Intel oneMKL for matrix-vector multiplication and oneMKL PARDISO linear solver to compute the action of $(H-E)^{-1}$ on vectors. The essence of this approach is to apply the standard Lanczos algorithm to the resolvent $(H-E)^{-1}$, where $E$ is a target energy, in our case selected randomly from the middle half of the spectrum.
Figure S4: Mean level spacing $\Delta\epsilon$ vs. system size $L$ for $\Delta=2.0$ and $\Delta_{2}=0.1$. Dark blue points correspond to ED data for $L=18,20,22$ and orange point to the extrapolated value for $L=24$, used in subsequent shift-invert calculations.
Using the resolvent of the Hamiltonian turns a dense part of the spectrum into a sparse one, facilitating the convergence of Lanczos procedure. In this way, we can sample the relevant eigenstates and approximate moments of $j_{nn}$ without performing a full diagonalization. We find $M=30$ Lanczos steps and $100-150$ states per momentum sector being enough for the results to converge. However, there is some caution required, as the action of the resolvent on a vector results in a system of linear equations $(H-E)|w\rangle=|v\rangle$, that is severely ill-conditioned, with condition number $\kappa\propto\exp{L}$. Hence, iterative solvers usually fail to reach satisfying precision, and direct solvers have to be used instead, limiting accessible systems sizes.
To compute the level sensitivity $R$ and spin conductivity $\sigma_{\text{CE}}$ (Eq. (3) and Eq. (6) in the main text), we also require the mean level spacing $\Delta\epsilon$ in the central part of the spectrum. Taking into account the exponential dependence of $\Delta\epsilon$ on the size of the system, it is straightforward to extrapolate this value from the ED data for $L=18,20,22$. An example of such an extrapolation for the case of $\Delta=2.0$ and $\Delta_{2}=0.1$ is shown in Fig. S4.
S0.3 Folded XXZ model and its properties as the toy model for the easy-axis regime
The ’folded’ XXZ model can be regarded as the easy-axis limit, $\Delta\to\infty$, of the XXZ chain. In order to calculate the matrix elements of the spin current operator, $j_{nn}$, we study this model in the presence of a nonzero flux $\varphi$,
$\tilde{H}=\frac{J}{2}\sum_{l}P_{l}[\leavevmode\nobreak\ \mathrm{e}^{-i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c.}\leavevmode\nobreak\ ],$ (S2)
Here, the projector, $P_{l}=[S^{z}_{l+2}+S^{z}_{l-1}]^{2}$, ensures that the spin exchange does not modify the total number of domain walls (DW). The GCE the lower bound, presented in the main text in Eq. (5) is valid also for the ’folded’ model, i.e. $\overline{j_{nn}^{2}}=1/4$ and appears to be tight in this limit (see Fig. 1). On the other hand, the CE results appear to be strictly zero, $\overline{j_{nn}^{2}}=0$, (at least numerically) for any finite $L$, for which we give further analytical arguments and explanation.
The aim is to present arguments that there are no diagonal current elements $j_{nn}$ in any eigenstate at $S^{z}_{tot}=0$, sticking to systems with even $L$. It is convenient to start from the reference antiferromagnetic (AFM) state and subsequently add additional domain walls $N_{W}=0,2,4\ldots$. While in the $N_{W}=0$ sector there is no dynamics within the model, Eq. (S2), the first nontrivial subspace is $N_{W}=2$.
Figure S5: Sketch showing dynamics of the ’folded’ model with $N_{W}=2$ domain walls (marked with dotted lines) in the antiferromagnetic background.
The pair of DW can move and behave as charged solitons (particles), see Fig. S5. For $N_{W}=2$ they can be described by the toy model,
$H_{\text{t}}=t\sum_{l}[\leavevmode\nobreak\ \mathrm{e}^{i\varphi}d^{\dagger}_{l+2}d_{l}+\mathrm{e}^{-i\varphi}c^{\dagger}_{l+2}c_{l}+\mathrm{H.c.}\leavevmode\nobreak\ ],$ (S3)
where $c_{l},d_{l}$ represent, respectively, positive and negative solitons (DW) on a bond between sites $l$ and $l+1$, and $t=J/2$. DW move on one sublattice (even or odd), but are impenetrable $d^{\dagger}_{l}d_{l}c^{\dagger}_{l}c_{l}=0$ and cannot cross each other. The two-particle problem can be solved exactly, but here the key feature is that the energies are independent of $\varphi$, hence $j_{nn}={\rm d}\epsilon_{n}(\varphi)/{\rm d}\varphi=0$. This can be confirmed in different ways. The eigenproblem for the translationally-invariant toy model with periodic boundary conditions can be written down in the general form,
$|\Psi\rangle=\sum_{m}f_{m}|\tilde{\psi}_{m}\rangle,\quad|\tilde{\psi}_{m}\rangle=\sum_{j}\mathrm{e}^{i2Kj}T_{2j}c^{\dagger}_{2m+1}d^{\dagger}_{1}|0\rangle,$ (S4)
where $K$ is a wave-vector and $T_{2j}$ is the translation operator that shifts positions of both particles by $2j$ sites. The latter translation takes into account that in the toy model, Eq. (S3), DW hops by two sites, that is, DW stays on the same sublattice. In the new basis the Hamiltonian acts as
$H_{\text{t}}|\tilde{\psi}_{1}\rangle=t\leavevmode\nobreak\ \mathrm{e}^{i\varphi}(1+\mathrm{e}^{i2K})|\tilde{\psi}_{2}\rangle\leavevmode\nobreak\ ,$
$H_{\text{t}}|\tilde{\psi}_{l}\rangle=t\leavevmode\nobreak\ \mathrm{e}^{-i\varphi}(1+\mathrm{e}^{-i2K})|\tilde{\psi}_{l-1}\rangle+t\leavevmode\nobreak\ \mathrm{e}^{i\varphi}(1+\mathrm{e}^{i2K})|\tilde{\psi}_{l+1}\rangle\leavevmode\nobreak\ ,$
$H_{\text{t}}|\tilde{\psi}_{L/2-1}\rangle=t\leavevmode\nobreak\ \mathrm{e}^{-i\varphi}(1+\mathrm{e}^{-i2K})|\tilde{\psi}_{L/2-2}\rangle\leavevmode\nobreak\ .$ (S5)
The resulting Hamiltonian matrix is tri-diagonal, where all non-diagonal elements are equal $b_{l}=t\leavevmode\nobreak\ \mathrm{e}^{i\varphi}(1+\mathrm{e}^{i2K})$. Most importantly, the eigenenergies, $\epsilon_{n}$, are determined by the products $B_{l}=b_{l}b^{*}_{l}$ that are real and do not involve $\varphi$. Hence, $\epsilon_{n}$ are independent of $\varphi$ and consequently $j_{nn}=0$.
In the same way, one can introduce additional DW, keeping $S^{z}_{tot}=0$. In particular, the case with $N_{W}=4$ domain walls is sketched in Fig. S6.
Figure S6: The same as in Fig. S5 but with $N_{W}=4$ domain walls.
In the above case we introduced solitons (DW) both on even and odd sites, and they can again move by hopping by two sites. They are clearly also other possibilities to generate $N_{W}>2$. It is important that positive and negative solitons are always created in pairs and that they are effectively impenetrable/noncrossing being the origin of $j_{nn}=0$, confirmed numerically for all configurations.
It is relevant to point out the analogy of physics of ’folded’ model and Eq. (S3) to the $t$ model [26], i.e., the Hubbard model at $U\to\infty$ with magnetic flux (acting oppositely on up/down spins),
$H_{\text{im}}=t\sum_{l}[\leavevmode\nobreak\ \mathrm{e}^{-i\varphi}c^{\dagger}_{l+1\uparrow}c_{l\uparrow}+\mathrm{e}^{i\varphi}c^{\dagger}_{l+1\downarrow}c_{l\downarrow}+\mathrm{H.c.}\leavevmode\nobreak\ ],$ (S6)
with imposed hard-core $U\to\infty$ repulsion $n_{\downarrow l}n_{\uparrow l}=0$. This model has the same properties as the ’folded’ model with respect to the spin current $j_{nn}={\rm d}\epsilon_{n}(\varphi)/{\rm d}\varphi=0$, provided that $N_{\uparrow}=N_{\downarrow}$ [26], whereby in Eq. (S6) the structure of (only nearest-neighbor) hopping is simpler.
Finally, the analysis of the ‘folded‘ model has the relevant consequence even for perturbed systems at $\Delta\to\infty$. Namely, the introduction of the next nearest neighbor interaction, $\Delta_{2}>0$, influences neither the impenetrability of DW nor the independence of energies on the flux, $\epsilon_{n}(\varphi)=\epsilon(0)$. Consequently one gets $j_{nn}=0$ also for the perturbed ‘folded‘ model. It explains why we also find numerically that $\sigma_{CE}$ decreases with $\Delta$ even at finite $\Delta_{2}$. On the other hand, $g=J_{2}$ perturbation directly allows for crossing of DW, hence even $\Delta\to\infty$ limit is non-trivial and the perturbed ‘folded‘ model can serve as the nontrivial starting point.
S0.4 Bound on the diagonal matrix elements of the spin current
In the main text, we introduced a product of two LIOMs, $Q=Q_{1}Q_{3}$, and used it to prove an explicit bound on the diagonal matrix elements $\overline{j^{2}_{nn}}$ in the easy-axis XXZ chain, see Eq. (5). In what follows, we argue that similar reasoning can be generalized for other translationally invariant, integrable models. Namely, we demonstrate that if the spin current has nonvanishing projection on any product of $m$ LIOMs,
$\langle jQ_{1}Q_{2}\ldots Q_{m}\rangle\neq 0,$ (S7)
then the Mazur-bound can be expressed in terms of a power-law dependence on $L$
$\overline{j^{2}_{nn}}\geq\frac{\langle jQ_{1}Q_{2}\ldots Q_{m}\rangle^{2}}{\langle Q^{2}_{1}Q^{2}_{2}\ldots Q^{2}_{m}\rangle}\geq\frac{\rm const}{L^{2m}}.$ (S8)
Here, $Q_{i}$, $i=1,\ldots,m$, denote arbitrary LIOMs. In order to demonstrate the inequality on the right-hand-side of Eq. (S8), we note that the normalization of $Q_{i}$ is not important. One may multiply any LIOM by a constant, $Q_{i}\to c_{i}Q_{i}$, and the constant $c_{i}$ cancels out in Eq. (S8). Therefore, we assume that $Q_{i}$ are translationally invariant and extensive operators, $Q_{i}=\sum_{l=1}^{L}(Q_{i})_{l}$. Here, $(Q_{i})_{l}$ is the local density of $Q_{i}$ supported on a few sites in the vicinity of site $l$. Then, the term in the numerator in Eq. (S8) reads
$\langle jQ_{1}...Q_{m}\rangle=\sum_{l_{0}=1}^{L}\sum_{l_{1}=1}^{L}\ldots\sum_{l_{m}=1}^{L}\langle(j)_{l_{0}}(Q_{1})_{l_{1}}\ldots(Q_{m})_{l_{m}}\rangle.$ (S9)
In the case of extensive operators, the corresponding densities $(Q_{i})_{l}$ and $(j)_{l}$ are independent of the system size. Therefore, all summands in Eq. (S9) are also $L$-independent and they are either zero or are of the order one ($\sim L^{0}$). Due to the assumption posed by Eq. (S7), some terms must be nonzero. If a certain term is nonzero then, due to translational invariance, there are $L$ identical summands in Eq. (S9). Consequently, the left-hand side of Eq. (S9) should be (at least) of the order $L$, unless there is some peculiar cancellation of positive and negative summands. However for the discussion of the level sensitivity, one may safely assume that $\langle jQ_{1}\ldots Q_{m}\rangle$ is of order $L$. The denominator in Eq. (S8) represents squared norm of the operator $Q_{1}\ldots Q_{m}$ so it is positive and nonzero. One may write down $\langle Q^{2}_{1}...Q^{2}_{m}\rangle$ in the form analogous to Eq. (S9). The resulting expression contains summations of terms which are either zero or are of order one and summations are carried out over $2m$ lattice indexes. We use the simplest bound $\langle Q^{2}_{1}\ldots Q^{2}_{m}\rangle\leq cL^{2m}$, where $c$ is a constant of the order one. Consequently, one obtains the bound from Eq. (S8), $\overline{j^{2}_{nn}}\geq{\rm const}/L^{2m}$. This combined together with the definition of the level sensitivity, see Eq. (3) in the main text, means that $R\geq{\rm const}/(L^{m+1}\Delta\epsilon)$. Since the level spacing, $\Delta\epsilon$, decays exponentially with $L$, the power-law term, $L^{m+1}$, may introduce only logarithmic corrections, $\sim[\log(\Delta\epsilon)]^{m+1}$, to the latter dependence. Consequently, Eq. (S7) is sufficient for the level sensitivity to exhibit a significant decrease upon introducing IBP from $R\sim 1/\Delta\epsilon$ (in the integrable case) to $R\sim 1/\sqrt{\Delta\epsilon}$ (in the chaotic system). We stress that the bound on the right-hand side of Eq. (S8) is not tight, however it is sufficient for the purpose of the present discussion.
S0.5 Nonzero magnetization sectors
In the main text, we have discussed results obtained for $S^{z}_{tot}=0$. In the case of integrable chain, we have found substantial differences between the level sensitivities obtained in the canonical ensemble ($R_{CE}$) and the grand canonical ensemble ($R_{GCE}$). This indicates that $R_{CE}$ for $S^{z}_{tot}=0$ must be very different from $R_{CE}$ in sectors with other magnetization, $m=S^{z}_{tot}/L$. However, one expects that both ensembles become equivalent at least in the chaotic regime.
Figure S7: Level sensitivity $R_{CE}$ vs. perturbation strength $\Delta_{2}$ within magnetization sectors $S^{z}_{tot}=0,1,2,3,4$, calculated via full ED in system with $L=20$, for (a) $\Delta=1.5$ and (b) $\Delta=2.0$. In panels (c) and (d) we show square of level sensitivity rescaled by the Hilbert space dimension and $m^{2}$, for $\Delta=1.5$ and $\Delta=2.0$ respectively. Dashed black line indicates a power-law dependence $\propto(\Delta_{2})^{-2}$.
In order to confirm both these expectations, we study the XXZ model [Eq. (1) in the main text] with perturbation $g=\Delta_{2}$ and $H^{\prime}=\sum_{l}S^{z}_{l+2}S^{z}_{l}$. Figures S7(a) and S7(b) show $R_{CE}$ as a functions of $\Delta_{2}$ for $L=20$ and $S^{z}_{tot}=0,1,2,3,4$. Here, $S^{z}_{tot}=1$ means that we applied one spin flip with respect to $S^{z}_{tot}=0$. One observes strong $S^{z}_{tot}$-dependence for $\Delta_{2}=0$ that diminishes when the integrability-breaking perturbation increases. For $\Delta>\Delta^{*}$, i.e. in the chaotic regime, $R_{CE}$ appears to change smoothly with magnetization. This smooth dependence is consistent with the equivalence of CE and GCE.
Finally, we demonstrate for the chaotic regime that the magnetization-dependence of $R_{CE}$ is consistent with $S^{z}_{tot}$-dependence of the spin conductivity. The latter consistency is set by the RMT-relation [Eq. (7) in the main text]. Spin conductivity, $\sigma$, has been studied numerically in Ref. [15], see Appendix A therein, for sectors with various magnetization, $m=S^{z}_{tot}/L$. The following relation has been established
$\sigma(m)$ $\simeq$ $\sigma(m=0)+{\rm const}\frac{m^{2}}{(\Delta_{2})^{2}}$ (S10)
$\simeq$ ${\rm const}\frac{m^{2}}{(\Delta_{2})^{2}},$ (S11)
and in Eq. (S11) we assume that $m/\Delta_{2}$ is sufficiently large. Following Eq. (7) in the main text, one expects that $R^{2}_{CE}\Delta\epsilon$ should exhibit the same dependence on $m$ and $\Delta_{2}$. As a rough estimate of the level spacing we take $\Delta\epsilon\propto 1/Z_{m}$, where $Z_{m}$ is the dimension of the subspace with magnetization $m$. Figures S7(c) and S7(d) show $R^{2}_{CE}/(m^{2}Z_{m})$ for $S^{z}_{tot}=1,2,3,4$. In the chaotic regime one observes nearly overlapping curves which decay as $1/(\Delta_{2})^{2}$. The deviations from this dependence show up, as expected, for small $m/\Delta_{2}$.